OpenStudy (anonymous):

9. A bag contains 5 white marbles, 4 black marbles, and 1 red marbles. Find the probability of picking a white marble? two-fifths one-half one-fifth one-tenth 10. In a bag of marbles you have 3 green, 2 yellow, 6 blue, and 9 red marbles. What is the probability of picking a red marble from the bag? three-twentieths one-tenth three-tenth nine-twentieths 11. In a bag of marbles you have 3 green, 2 yellow, 6 blue, and 9 red marbles. What is the probability of picking a yellow marble from the bag? three-twentieths nine-twentieths o

OpenStudy (igreen):

Q: 9 There are 10 marbles total. You want to know the probability of choosing white. There are 5 marbles total. 5/10 is the probability of picking a white marble. We can simplify this fraction to 1/2

OpenStudy (igreen):

Q: 10 There are a total of 20 marbles. There are 9 red marbles. So the probability of picking a red marble is 9/20.

OpenStudy (igreen):

Q: 11 There are 20 marbles total There are 2 yellow marbles. So the probability of picking a yellow marble is 2/20 or 1/10.

OpenStudy (igreen):

@chappy0729

OpenStudy (anonymous):

thankn you so much !!!!!! @iGreen

OpenStudy (igreen):

No problem.

OpenStudy (anonymous):

can you help me one more? @iGreen

OpenStudy (igreen):

Sure

OpenStudy (anonymous):

14. A bag has 7 red marbles, 8 blue marbles and 3 black marbles. What is the probability of picking a blue marble, replacing it, and then picking a red marble? seven one-hundred-sevenths one-ninth two twenty-sevenths fourteen eighty-firsts 15. A bag has 2 yellow, 7 red, and 1 green marble. What is the probability of picking a red marble, replacing it, and then picking another red marble? forty-nine hundredths one fiftieth seven hundredths one twenty-fifth 16. A bag has 2 yellow, 7 red, and 1 green marble. What is the probability of picking a green marble, replacing it, and then picking a yellow marble? seven hundredths one fiftieth forty-nine hundredths one twenty-fifth 17. A bag has 2 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a green marble, do not replace it, and pick a red marble? one twenty-eighth three twenty-eighths nine fifty-sixths one-eighth 18. A bag has 2 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a red marble, do not replace it, and pick yellow marble? three twenty-eighths one twenty-eighth nine fifty-sixths one-eighth 19. A bag has 2 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a yellow marble, do not replace it, and pick another yellow marble? nine fifty-sixths three twenty-eighths one twenty-eighth one-eighth

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

A bag has 7 red marbles, 8 blue marbles and 3 black marbles. What is the probability of picking a blue marble, replacing it, and then picking a red marble? There are 18 marbles total. The probability of taking a blue marble is 8/18 or 4/9. The probability of taking a red is 7/18. Multiply these two to find the probability. 4/9 * 7/18 = 14/81. So the probability of picking a blue and then a red is 14/81.

OpenStudy (igreen):

A bag has 2 yellow, 7 red, and 1 green marble. What is the probability of picking a red marble, replacing it, and then picking another red marble? There are 10 marbles total. The probability of taking a red marble is 7/10. Multiply this by 7/10 to find the probability of picking a red and then another red. 7/10 * 7/10 = 49/100. So the probability of picking a red and then another red is 49/100.

OpenStudy (igreen):

A bag has 2 yellow, 7 red, and 1 green marble. What is the probability of picking a green marble, replacing it, and then picking a yellow marble? The probability of picking a green marble is 1/10. The probability of picking a yellow marble is 2/10. Multiply these to find the probability of picking a green marble and then a yellow marble. 1/10 * 2/10 = 1/50. So the probability of picking a green marble and then a yellow marble is 1/50.

OpenStudy (igreen):

A bag has 2 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a green marble, do not replace it, and pick a red marble? The probability of picking a green marble is 2/8 or 1/4. The probability of picking a red marble AFTER you have picked a green marble is 3/7. Multiply these two to find the probability. 1/4 * 3/7 = 3/28. So the proabability of picking a green marble and then a red marble is 3/28.

OpenStudy (igreen):

A bag has 2 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a red marble, do not replace it, and pick yellow marble? The probability of picking a red marble is 3/8. The probability of picking a yellow marble after the red marble is 3/7. Multiply these to find the probability. 3/8 * 3/7 = 9/56. So the probability of picking a red marble and then a yellow marble is 9/56.

OpenStudy (igreen):

A bag has 2 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a yellow marble, do not replace it, and pick another yellow marble? The probability of picking a yellow marble is 3/8. The probability of picking a yellow marble after you already picked a yellow one is 2/7. Multiply these to find the probability. 3/8 * 2/7 = 3/28. So the probability of picking a yellow marble and then another is 3/28.

OpenStudy (igreen):

@chappy0729

OpenStudy (anonymous):

Thank you so much. @iGreen

OpenStudy (igreen):

No problem.